A shopkeeper bought 300 eggs at 80 paisa each. 30 eggs were broken in transport and then he sold the remaining eggs at one rupees .find his gain or loss per cent

Grade:8

1 Answers

Debarati
2 years ago
CP of eggs= 300×80p= 24000p= Rs. 240
No. of broken eggs= 30
No. of eggs remaining= 300-30= 270
SP of eggs= 270×1= Rs. 270
Clearly, the shpkeeper has gained from this transaction since 270>240, ie, SP>CP.
Now, gain= 270-240= Rs. 30
Gain %= (Gain/CP)×100= (30/240)×100= 12.5%
\thereforeGain %= 12.5%

I hope you understood.